# Network Modelling of Asynchronous Change-Points in Multivariate Time Series

### v1

@@ Abstract
This article introduces a novel Bayesian method for asynchronous change-point detection in multivariate time series. This method allows for change-points to occur earlier in some series (leading series) followed, after a short delay, by change-points in some other series (lagging series). Such dynamic dependence structure is common in fields such as seismology and neurology where a latent event such as an earthquake or seizure causes certain sensors to register change-points before others. We model these lead-lag dependencies via a latent directed graph and provide a hierarchical prior for learning the graph’s structure and parameters. Posterior inference is made tractable by modifying particle MCMC methods designed for univariate change-point problems. We apply our method to both simulated and real datasets from the fields of seismology and neurology. In the simulated data, we find that our method outperforms competing Bayesian methods in settings where the change-point locations are dependent across series. In the real data applications we show that our model can also uncover interpretable network structure.
Keywords: Multivariate, Change-Point, Asynchronous, MCMC, Particle Gibbs, Network, Time Series
@@ 1 Introduction
Change-point models are employed to infer the locations of abrupt changes in the distribution of a time series. These models assume that some combination of the distribution’s parameters shift at the change locations. Change-point analysis plays a critical role in modelling non-stationary systems, where structural changes arise from external shocks or internal dynamics. A model that correctly detects such changes leads to enhanced predictive performance and supports more reliable decision-making by accounting for shifts in underlying processes. The accurate detection of change-points also offers insight into the causal factors driving the behaviour of a dynamic system. Change-point models have been applied in many fields such as medicine (Liu et al.,, 2017; Kass-Hout et al.,, 2012), finance and economics (Chib,, 1998; Bauwens et al.,, 2017; Maheu and Song,, 2018) and hydrology (Alaya et al.,, 2020).
When analysing multiple time series, it is common to observe patterns such as simultaneous changes or short delays in changes across subsets of dependent series. This is often due to a common latent process driving the change-points across different series. A particularly important pattern, observed in many applications is the lead-lag relationship, where leading series tend to experience change-points shortly before their corresponding lagging series. Identifying lead-lag relationships is important both for prediction — since a change in a leading series may signal an imminent change in a lagging series — and for inference, by providing insight into the latent processes connecting the series.
In this article we present a novel Bayesian method for modelling multivariate change-points in a variety of settings. Our approach encourages change-points to appear earlier in some (leading) series, followed shortly after by change-points in other (lagging) series. To encode these lead-lag relationships we use a latent directed graph where vertices represent the series and a directed edge from one vertex/series to another indicates that one series leads the other. The strength of the lead-lag relationships and the delays between change-points are controlled via a set of edge parameters. We assume that both the edge set and edge parameters are unknown and will be inferred in addition to the change-point locations. To perform posterior inference we extend the efficient particle Gibbs sampler of Whiteley et al., (2011) used in univariate change-point models to the multivariate setting.
@@
Although lead-lag relationships arise in many fields, our work is motivated by two specific examples. The first concerns the detection of seismic events using observations from multiple ground motion sensors. In this setting, an event such as an earthquake induces abrupt changes in the amplitude and frequency content of the sensor signals as the primary (P) and secondary (S) waves arrive. Sensors located closer to the fault line typically register change-points earlier than those further away, giving rise to a lead-lag relationship across the multiple series. Here, the lead-lag relationships are driven by spatial proximity, and modelling them may improve detection accuracy of weaker events at more distant sensors.
@@
The other motivating example is the detection of seizures using electroencephalogram (EEG) signals. EEG analysis involves recording the brain’s electrical activity over time using electrodes placed on the scalp. The occurrence of a seizure results in sudden changes in the amplitude and rhythmic patterns of the EEG signals. These changes are initially detected in the electrodes closest to the seizure onset zone and typically propagate to nearby electrodes shortly thereafter as the seizure spreads to other regions of the brain. In this context, lead-lag relationships are driven by the seizure’s dynamics and propagation pathway. Inferring this pathway is crucial for understanding the brain’s network structure and for developing personalized treatment strategies aimed at preventing the seizure from spreading.
The literature on performing change-point detection in the univariate setting is quite large. Frequentist approaches date back to Page, (1954) and generally involve minimising cost functions or recursively applying tests to detect multiple changes. Recent advances in this area include the Pruned Exact Liner Time (PELT) method (Killick et al.,, 2012) and wild binary segmentation (Fryzlewicz,, 2014). For a selective review of these techniques, see for example Truong et al., (2020).
Univariate Bayesian methods can be traced back to Chernoff and Zacks, (1964) and require a prior distribution for the change-point locations. These models typically assume the change-points follow a Markov chain, differing in how they define the transitions and whether the number of changes is known or unknown. When the number of changes is known, we may use a finite state hidden Markov model (HMM), see for example Chib, (1996). For an unknown number of changes, we may compare various finite state models with Bayes factors (Chib,, 1998), however, this quickly becomes intractable when comparing many models. An alternative solution is to use the infinite hidden Markov model of Beal et al., (2002), a nonparametric extension of Chib, (1996) which allows for a countably infinite number of states.
An alternative and widely used approach to modelling an unknown number of change-points is to use Product Partition Models (PPMs) (Hartigan,, 1990), which generalise several previous methods, such as that of Yao, (1984). These models specify a distribution over the random partition induced by the change-points and can also be formulated as a Markov chain (Barry and Hartigan,, 1992). However, unlike HMMs which may re-enter previous states, a change-point under a PPM always introduces a completely new state. Recent work employing PPMs include (Nyamundanda et al.,, 2015; García and and,, 2019; Corradin et al.,, 2022). While conceptually simple to specify, Bayesian methods often require Monte Carlo sampling in order to obtain posterior estimates, see for example Fearnhead, (2006).
Less attention has been given to the multivariate setting where changes may affect subsets of the series at synchronised or nearly synchronised times. Motivated by applications in genomics, much of the frequentist literature focuses on detecting synchronized changes in the mean of Gaussian random variables. When changes are assumed to affect all series simultaneously, Zhang et al., (2010) propose a scanning procedure which sums the chi-squared statistics across each series and show that their method leads to improved detection accuracy. A more general approach was introduced by Siegmund et al., (2011) which includes the method of Zhang et al., (2010) as a special case.
When synchronised changes are observed on only a small subset of the series, Jeng et al., (2013) proposed the Proportion Adaptive Segment Selection (PASS) procedure, which automatically adapts to the unknown number of series which experience synchronised change-points. An alternative method is proposed by Wang and Samworth, (2018) who first produce a sparse projection of the series before applying the cumulative sum (CUSUM) test to detect a single change in the resulting series. Multiple change-points can then be handled by combining this with the binary segmentation method of Fryzlewicz, (2014).
Similar to the frequentist literature, most existing Bayesian methods are designed for analysing synchronised change-points and build upon existing univariate methods. For example, Fan and Mackey, (2017) proposed an empirical Bayes method, such that for each fixed time, the prior change-point probabilities are the same across all series, and can vary with time. A fully Bayesian method, motivated by the univariate PPM approach (Barry and Hartigan,, 1992), was proposed by Quinlan et al., (2024). This method also encourages synchronized changes by specifying a joint distribution for the vector of prior change-point probabilities for each fixed time. A related problem was considered by Bardwell and Fearnhead, (2017) who proposed the Bayesian Abnormal Segments Detector (BARD) method. This method models the change-point locations with a Markov chain and attempts to identify segments where a subset of the series differs from some baseline behaviour. The three methods of Fan and Mackey, (2017); Bardwell and Fearnhead, (2017); Quinlan et al., (2024) encourage subsets of the series to exhibit synchronised changes a priori, but ignore any dependence which may arise when change-points occur asynchronously. Similarly, they provide no method for inferring which subsets of the series exhibit dependence (other than manually inspecting the posterior change-point probabilities).
A Bayesian method that allows for both synchronous and asynchronous dependence in the change locations was proposed by Hallgren et al., (2024) who assumed that the series are connected via a known weighted undirected graph. Under this model, series which are connected on the graph are more likely to experience changes at similar, but not necessarily synchronised times. While this method allows for asynchronous dependence, it does not account for lead-lag relationships and requires us to supply a graph linking the series which, in general, may not be known.
@@
Our approach to modelling multivariate change-points is similar to Hallgren et al., (2024). We too assume that the series are connected via a latent weighted graph. However, unlike the approach of Hallgren et al., (2024), we do not assume the graph is known a priori and we provide a hierarchical prior to enable learning on the graph parameters. Since we are interested in lead-lag relationships we use a directed graph to account for directed dependence.
@@
The remainder of this article is organised as follows, Section 2 introduces the proposed model and discusses its properties. In Section 3 we describe a novel blocked particle Gibbs scheme to obtain posterior samples. In Section 4 we demonstrate our methods through a comparative simulation study and two real data examples before concluding with a discussion in Section 5.
@@ 2 The Model
In this section we introduce our modelling approach, which we refer to as Network Change-Point (NetCP) model, and related notation. We consider a multivariate time series of dimension $d$, observed at times $t=1,...,T$. Let $\boldsymbol{Y}_{t}=(Y_{1,t},...,Y_{d,t})^{{}^{\prime}}$ denote the vector of observations at time $t$ and $\boldsymbol{Y}=(\boldsymbol{Y}_{1},...,\boldsymbol{Y}_{T})$ be the $(d\times T)$ matrix which holds the observations. Conditional on a set of change-point locations, the data in each series is partitioned into non-overlapping segments such that the change-points correspond to the segment end-points. We use the notation $\boldsymbol{Y}_{j,s:t}=(Y_{j,s+1},...,Y_{j,t})$ to denote the segment of observations in series $j$ between change-points $s$ and $t$.
We model the change-point locations with a multivariate hidden state process. Each observation $Y_{j,t}$ is assigned a hidden state $X_{j,t}$, which denotes the time of the most recent change-point in series $j$ before time $t$. Our prior distribution over the hidden states $\boldsymbol{X}$ is parameterised according to a latent directed graph linking the pairs of series that are dependent. We focus on posterior inference for both the hidden states and the graph parameters.
@@ 2.1 Likelihood Model
We first describe the likelihood model conditional on the hidden state process. Recall that $X_{j,t}$ denotes the time of the most recent change-point in series $j$ prior to time $t$, then event $X_{j,t}=t-1$ indicates that a change-point occurred at time $t-1$. So, conditional on $\boldsymbol{X}=\boldsymbol{x}$, the set of change-point locations in series $j$ is given by
$\boldsymbol{\tau}_{j}=\{i\in\{0,...,T-1\}:x_{j,i+1}=i\}\cup\{T\}$
The change-points partition the observations in each series $j$ into $|\boldsymbol{\tau}_{j}|-1$ segments, where the $k^{th}$ segment is given by $\tau_{j,k}:\tau_{j,k+1}=\{\tau_{j,k}+1,...,\tau_{j,k+1}\}$. We assume that the observations in the $k^{th}$ segment have density $f_{j}(\boldsymbol{y}_{j,\tau_{j,k}:\tau_{j,k+1}}|\theta_{j,k})$ where the parameter $\theta_{j,k}\in\Theta$ is segment specific and is drawn independently from the density $f_{j}(\cdot)$.
For computational reasons, (see Barry and Hartigan,, 1993; Fearnhead,, 2006), the densities $f_{j}(\cdot|\theta)$ and $f_{j}(\cdot)$ commonly chosen to be conjugate, enabling us to analytically marginalise over the segment specific parameters:
[…]
The resulting density $f_{j}(\boldsymbol{y}_{j,s:t})$ is referred to as the marginal segment density. Throughout this paper we will assume that (1) is available analytically, or can be approximated numerically in low dimensions. Under these assumptions, the joint likelihood may be expressed as
@@
$f_{j}(y_{j,t}|x_{j,t},\boldsymbol{y}_{j,0:t-1})=\frac{f_{j}(\boldsymbol{y}_{j,x_{j,t}:t})}{f_{j}(\boldsymbol{y}_{j,x_{j,t}:t-1})}$ (4)
can be computed from the marginal segment densities given in Equation (1). The likelihood structure given in Equations (1) - (4) has been widely used in both the univariate and multivariate change-point literature (Barry and Hartigan,, 1993; Fan and Mackey,, 2017; Hallgren et al.,, 2024).
The specific choice of $f_{j}(\cdot|\theta)$ and $f_{j}(\cdot)$ depends on the application area. In Section 4 we apply our model to seismology and EEG data, where the focus is detecting changes in the variability and/or autocorrelation structure of each signal. To do this, this we employ an independent autoregressive model for the data within each segment of each series, such that both the lag coefficients and variance can vary across segments. In particular, we assume the segment specific parameters are $\theta:=(\boldsymbol{\phi},\sigma^{2})$ where $\boldsymbol{\phi}\in\mathbb{R}^{L}$ is the vector of autoregressive parameters and $L$ is the number of lags. For a given segment $\boldsymbol{y}_{j,s:t}=\{y_{j,s+1},...,y_{j,t}\}$, the likelihood is then given by
@@
where $\mathcal{N}_{t-s}(\cdot|\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the density of the $t-s$ dimensional normal distribution, $\boldsymbol{I}_{t-s}$ denotes the $(t-s)\times(t-s)$ identity matrix and
$\boldsymbol{H}_{j,st}=\begin{bmatrix}y_{j,s}&y_{j,s-1}&\cdots&y_{j,s-L+1}\\ y_{j,s-1}&y_{j,s-2}&\cdots&y_{j,s-L}\\ \vdots&\vdots&\ddots&\vdots\\ y_{j,t-1}&y_{j,t-2}&\cdots&y_{j,t-L}\end{bmatrix}$ (6)
is a $(t-s)\times L$ matrix with rows containing the lagged values for each observation in the segment. We place a conjugate normal-inverse-gamma prior over $\theta=(\boldsymbol{\phi},\sigma^{2})$:
@@
where $\mathcal{IG}(\sigma^{2}|\alpha,\beta)$ denotes the density of the inverse gamma distribution and $(\alpha_{j},\beta_{j},\delta_{j,1},...,\delta_{j,L})$ are a set of hyperparameters for series $j$. Integrating over $(\boldsymbol{\phi},\sigma^{2})$ within a given segment, we obtain the marginal segment density:
@@
where $\Gamma(\cdot)$ denotes the gamma function, $\boldsymbol{D}_{j}=\text{diag}(\delta_{j,1},...,\delta_{j,L})$, $\alpha_{j,st}=\alpha_{j}+(t-s)/2$, $\beta_{j,st}=\beta_{j}+\frac{1}{2}[\|\boldsymbol{Y}_{j,s:t}\|_{2}^{2}-\boldsymbol{E}_{j,st}\boldsymbol{D}_{j,st}\boldsymbol{E}_{j,st}^{T}]$, $\boldsymbol{E}_{j,st}=\boldsymbol{Y}_{j,s:t}\boldsymbol{H}_{j,st}$ and $\boldsymbol{D}_{j,st}=(\boldsymbol{H}_{j,st}^{{}^{T}}\boldsymbol{H}_{j,st}+\boldsymbol{D}_{j}^{-1})^{-1}$. This likelihood model was previously considered by Fearnhead and Liu, (2007) in the univariate setting and may be used to model data with a constant mean but which exhibits change-points in the variability and/or autocorrelation structure.
In practice, it is essential to carefully select the hyperparameters $(\alpha_{j},\beta_{j},\delta_{j,1},\ldots,\delta_{j,L})$ for each series to minimise the error rate in the change-point locations. A fully Bayesian approach involves placing priors on these parameters and updating them during posterior sampling. However, fixing the parameters can yield substantial computational savings, as each marginal segment density must only be computed once and can be reused. In the examples presented in Section 4, we present sensible default values.

### v2

# Network Modeling of Asynchronous Change-Points in Multivariate Time Series

@@ Abstract
This article introduces a novel Bayesian method for asynchronous change-point detection in multivariate time series. This method allows for change-points to occur earlier in some (leading) series followed, after a short delay, by change-points in some other (lagging) series. Such dynamic dependence structure is common in fields such as seismology and neurology where a latent event such as an earthquake or seizure causes certain sensors to register change-points before others. We model these lead-lag dependencies via a latent directed graph and provide a hierarchical prior for learning the graph’s structure and parameters. Posterior inference is made tractable by modifying particle MCMC methods designed for univariate change-point problems. We apply our method to both simulated and real datasets from the fields of seismology and neurology. In the simulated data, we find that our method outperforms competing methods in settings where the change-point locations are dependent across series. In the real data applications we show that our model can also uncover interpretable network structure.
Keywords: Segmentation; Particle Gibbs; Electroencephalogram; Lead-lag dependence
@@ 1 Introduction
Change-point models are used to infer the locations of abrupt changes in the distribution of a time series. These models assume that a potentially unknown number of change-points partition the data into a set of segments. The data within each segment are assumed to be generated from a common model, which varies from one segment to another. Examples include Gaussian models with changing means and variances [53, 73] or regression curves with discontinuous jumps [61]. These models are applied to a range of problems, including the detection of epileptic seizures [85], the identification of economic regime shifts [78], and the detection of earthquakes [91].
In this article, we present a novel Bayesian method for detecting change-points in multivariate time series. Our method allows for dependence between pairs of series such that the change-points in some (lagging) series are likely to occur shortly after those in other (leading) series. Identifying such lead-lag relationships is important both for prediction, since detecting a change in a leading series may signal an imminent change in a lagging series, and for inference, by providing insight into the latent processes connecting the series. We represent each series as a node on a latent directed graph whereby a directed edge from one series/node to another indicates the presence of a lead-lag relationship, the strength of which is controlled by a set of edge parameters.
@@
Although lead-lag relationships arise in many fields, this work is motivated by two specific examples. The first concerns the detection of seismic events using observations from multiple ground motion sensors. In this setting, an event such as an earthquake induces abrupt changes in the amplitude and frequency content of the sensor signals as the primary (P) and secondary (S) waves arrive. Sensors located closer to the fault line typically register change-points earlier than those further away, giving rise to a lead-lag relationship across the multiple series. Here, the lead-lag relationships are driven by spatial proximity, and modelling them may improve detection accuracy of weaker events at more distant sensors.
@@
The other motivating example is the detection of seizures using electroencephalogram (EEG) signals. EEG analysis involves recording the brain’s electrical activity over time using electrodes placed on the scalp. The occurrence of a seizure results in sudden changes in the amplitude and rhythmic patterns of the EEG signals. These changes are initially detected in the electrodes closest to the seizure onset zone and typically propagate to nearby electrodes shortly thereafter, as the seizure spreads to other regions of the brain. In this context, lead-lag relationships are driven by the seizure’s dynamics and propagation pathway. Inferring this pathway is crucial for understanding the brain’s network structure and for developing personalized treatment strategies aimed at preventing the seizure from spreading.
Early work on change-point detection took a frequentist approach. Univariate methods can be traced back to [81] who proposed the Cumulative Sum (CUSUM) test for detecting a single change-point in the mean. Multiple change-points can be handled by applying such tests recursively using algorithms such as Binary Segmentation (BS) [88] or Wild Binary Segmentation (WBS) [66], a stochastic extension to BS which is more sensitive to multiple changes. Another popular approach is to minimize a cost function which assesses the fit of a particular segmentation relative to the number of change-points. Recent work in this area includes the Pruned Exact Linear Time (PELT) method [76] which performs this minimization in linear time. Reviews of various methods in this area can be found in [72, 87].
In the multivariate setting, many frequentist approaches have focused on situations where change-points affect subsets of the series at synchronized times [[, see for example]]Jeng2013, Wang2018 and are therefore not appropriate for the problem at hand. Two exceptions to this are the methods of [91] and [65] who allow for change-points to occur with either a fixed delay or within some time window of those in other series. However, assuming a fixed delay may be overly restrictive, and these methods do not account for the lead-lag dependencies considered in this paper.
Bayesian approaches require a prior on the change-points and provide posteriors that quantify uncertainty about their number and locations. In the univariate setting, the early work by [53] and [92] proposed using a geometric renewal process, whereby the random arrival times determine the change locations. [49, 50] extended this by using Product Partition Models (PPMs) [71]. PPMs assume that the prior over the induced partition can be written as a product of segment-wise cohesion functions, which includes renewal processes as a special case (when the cohesions depend only on the segment length). Alternative approaches based on Markov chains have also been proposed. For instance, [54] models the change-points using a finite-state, unidirectional Markov chain and compares models with different numbers of change-points via Bayes factors. An overview of Bayesian methods and their computational aspects can be found in [58].
When considering multivariate time series, one could apply existing univariate Bayesian methods directly. For example, we may assume a single change-point prior which affects all series simultaneously [57] or assign independent change-point priors for different parameters of the model [83, 82]. However, these approaches are overly restrictive, assuming either complete dependence or independence in the change locations.
An alternative and more flexible approach is to allow each time point in each series to have its own change-point probability. Then induce dependence by assigning a joint prior over the vector of probabilities for each time. This was the approach of [70] who specified a Dirichlet distribution over all possible change-point configurations at each time. However, the number of parameters in the prior grows exponentially with the number of series, quickly becoming infeasible. An alternative, empirical Bayesian approach was proposed by [60] who assumed that the change-point probabilities are shared across series for each fixed time but may vary across time. This prior therefore encourages synchronized change-points across series.
Another fully Bayesian approach, inspired by the univariate PPM [49], is the method of [84]. Here, the vector of change probabilities at each time are generated by taking the logistic transform of random variables sampled from a multivariate Student’s t-distribution. Thus, the correlation between the change-point probabilities at each fixed time can be controlled through the covariance matrix. A related problem was considered by [48] who proposed the Bayesian Abnormal Segments Detector (BARD) method. This method models the change-point locations with a Markov chain and attempts to identify segments where a subset of the series differ from some baseline behavior. Although, this approach differs from standard change-point methods in that it assumes the series will return to some baseline behavior.
The multivariate approaches discussed so far rely on subsets of series changing simultaneously, making them unsuitable for the lead-lag dependencies considered here. A Bayesian method that allows for asynchronous changes was proposed by [69], who assumed that the series are connected via a known undirected graph. Under this method, series which are connected on the graph are encouraged to experience changes at similar times. However, because the graph is undirected, this method cannot capture directed lead–lag relationships. It also requires a prespecified graph, which may not be known a priori.
@@
Our approach to modeling multivariate change-points is similar to that of [69]. We also assume that the series are connected via a latent weighted graph. However, unlike their approach, we do not assume that the graph is known a priori, and we provide a hierarchical prior to enable learning on the graph parameters. Since we are interested in lead-lag relationships, we use a directed graph to account for directed dependence between pairs of series. We show how posterior inference can be performed by adapting the efficient particle Gibbs sampler of [90] to the multivariate setting.
@@
The remainder of this article is organized as follows: Section 2 introduces the proposed model and discusses its properties. In Section 3, we describe a novel blocked particle Gibbs scheme to obtain posterior samples. In Section 4, we demonstrate our method through a comparative simulation study and two real data examples before concluding with a discussion in Section 5. Proofs of results and background information on the particle Gibbs sampler are given in the Appendices. All code and data required to reproduce the plots, tables and analysis is available at github.com/carsonmckee/NetCP.
@@ 2 The Model
In this section, we introduce our modeling approach, which we refer to as the Network Change-Point (NetCP) model, and related notation. We consider a multivariate time series of dimension $d$, observed at times t=1,,Tt=1,...,T. Let $\boldsymbol{Y}_{t}=(Y_{1,t},...,Y_{d,t})^{{}^{\prime}}$ denote the vector of observations at time $t$ and $\boldsymbol{Y}=(\boldsymbol{Y}_{1},...,\boldsymbol{Y}_{T})$ be the $(d\times T)$ matrix which holds the observations. Conditional on a set of change-point locations, the data in each series is partitioned into non-overlapping segments such that the change-points correspond to the segment end-points. We use the notation $\boldsymbol{Y}_{j,s:t}=(Y_{j,s+1},...,Y_{j,t})$ to denote the segment of observations in series $j$ between change-points $s$ and $t$.
We model the change-point locations with a multivariate hidden state process, which we refer to as the NetCP process. Each observation $Y_{j,t}$ is assigned a hidden state $X_{j,t}$, which denotes the time since the last change-point before time $t$ in series $j$. That is, $X_{j,t}$ is the run length at time $t$ in series $j$. Therefore, at time $t$, the run length $X_{j,t}$ either resets to 1 (indicating a change-point) or increases by 1 (no change-point). Our prior over the hidden states is parameterised according to a latent directed graph, with edges linking the pairs of series that are dependent. We focus on posterior inference for both the hidden states and the graph parameters.
@@ 2.1 Likelihood Model
We first describe the likelihood model conditional on the hidden state process. Recall that $X_{j,t}$ denotes the run length at time $t$ in series $j$. Conditional on $\boldsymbol{X}=\boldsymbol{x}$, the set of change-point locations in series $j$ is given by
$\boldsymbol{\tau}_{j}=\{i\in\{0,...,T-1\}:x_{j,i+1}=1\}\cup\{T\}$
The change-points partition the observations in each series $j$ into $|\boldsymbol{\tau}_{j}|-1$ segments, where the $k$th segment is given by $\tau_{j,k}:\tau_{j,k+1}=\{\tau_{j,k}+1,...,\tau_{j,k+1}\}$. We assume that the observations in the $k$th segment of the $j$th series have density $f_{j}(\boldsymbol{y}_{j,\tau_{j,k}:\tau_{j,k+1}}|\theta_{j,k})$ where the parameter $\theta_{j,k}\in\Theta$ is segment specific and drawn independently from the density $f_{j}(\cdot)$.
For computational reasons, [[, see]]Barry1993, Fearnhead2006, the densities $f_{j}(\cdot|\theta)$ and $f_{j}(\cdot)$ are chosen to be conjugate, allowing for the analytic marginalization over the segment specific parameters, i.e.
$f_{j}(\boldsymbol{y}_{j,s:t})=\int_{\Theta}f_{j}(\boldsymbol{y}_{j,s:t}|\theta)f_{j}(\theta)d\theta$ (1)
where the density $f_{j}(\boldsymbol{y}_{j,s:t})$ is referred to as the marginal segment density. Throughout this paper, we assume that Equation (1) is available analytically or can be approximated numerically in low dimensions. Under these assumptions, the joint likelihood can be expressed as
@@
$f_{j}(y_{j,t}|x_{j,t},\boldsymbol{y}_{j,0:t-1})=\frac{f_{j}(\boldsymbol{y}_{j,(t-x_{j,t}):t})}{f_{j}(\boldsymbol{y}_{j,(t-x_{j,t}):t-1})}$ (4)
can be computed from the marginal segment densities given in Equation (1). The likelihood structure given in Equations (1) - (4) has been widely used in both the univariate and multivariate change-point literature [50, 60, 69].
The choice of $f_{j}(\cdot|\theta)$ and $f_{j}(\cdot)$ is problem-specific. In Section 4 we apply our model to seismology and EEG data, where the focus is to detect changes in the variability and/or autocorrelation structure of signals with constant mean. To do this, we use an independent autoregressive model for the data within each segment of each series, such that both the lag coefficients and variance can vary across segments. In particular, we assume that the segment-specific parameters are $\theta:=(\boldsymbol{\phi},\sigma^{2})$ where $\boldsymbol{\phi}\in\mathbb{R}^{L}$ is the vector of autoregressive parameters and $L$ is the number of lags. For a given segment $\boldsymbol{y}_{j,s:t}=\{y_{j,s+1},...,y_{j,t}\}$, the likelihood is given by
@@
where $\mathcal{N}_{t-s}(\cdot|\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the density of the tst-s dimensional normal distribution, $\boldsymbol{I}_{t-s}$ denotes the $(t-s)\times(t-s)$ identity matrix and
$\boldsymbol{H}_{j,st}=\begin{bmatrix}y_{j,s}&y_{j,s-1}&\cdots&y_{j,s-L+1}\\ y_{j,s+1}&y_{j,s}&\cdots&y_{j,s-L+2}\\ \vdots&\vdots&\ddots&\vdots\\ y_{j,t-1}&y_{j,t-2}&\cdots&y_{j,t-L}\end{bmatrix}$ (6)
is a $(t-s)\times L$ matrix with rows containing the lagged values for each observation in the segment. We place a conjugate normal-inverse-gamma prior over $\theta=(\boldsymbol{\phi},\sigma^{2})$ within each series:
@@
where $\mathcal{IG}(\sigma^{2}|\alpha,\beta)$ denotes the density of the inverse gamma distribution and $(\alpha_{j},\beta_{j},\delta_{j,1},...,\delta_{j,L})$ are the set of hyperparameters for series $j$. Integrating over $(\boldsymbol{\phi},\sigma^{2})$ within a given segment, we obtain the marginal segment density:
@@
where $\Gamma(\cdot)$ denotes the gamma function, $\boldsymbol{D}_{j}=\text{diag}(\delta_{j,1},...,\delta_{j,L})$, $\alpha_{j,st}=\alpha_{j}+(t-s)/2$, $\beta_{j,st}=\beta_{j}+\frac{1}{2}[\|\boldsymbol{Y}_{j,s:t}\|_{2}^{2}-\boldsymbol{E}_{j,st}\boldsymbol{D}_{j,st}\boldsymbol{E}_{j,st}^{{}^{\prime}}]$, $\boldsymbol{E}_{j,st}=\boldsymbol{Y}_{j,s:t}\boldsymbol{H}_{j,st}$ and $\boldsymbol{D}_{j,st}=(\boldsymbol{H}_{j,st}^{{}^{\prime}}\boldsymbol{H}_{j,st}+\boldsymbol{D}_{j}^{-1})^{-1}$. This likelihood model was previously considered by [64] in the univariate setting and is used to model data with a constant mean and change-points in the variability and/or autocorrelation structure.
To minimize the error rate in the change-point locations, it is essential to carefully select the hyperparameters $(\alpha_{j},\beta_{j},\delta_{j,1},\ldots,\delta_{j,L})$ for each series. A fully Bayesian approach involves placing priors on the hyperparameters and updating them during posterior sampling. However, fixing the parameters can yield substantial computational savings, as each marginal segment density must only be computed once and can be reused. In the examples presented in Section 4, we present sensible default values.